OpenStudy (anonymous):

solve for v v-4/5= 4 1/3

OpenStudy (anonymous):

the r.h.s is it..41/3 or is it a mixed no/??

OpenStudy (anonymous):

yes 4 1/3

OpenStudy (anonymous):

217/15 if not..a mixed no.!!

OpenStudy (anonymous):

sry its 4 then 1/3

OpenStudy (anonymous):

oh..wel its a mixed no. then..so..its equivalent is 4+1/3=13/3

OpenStudy (anonymous):

its a mixed number v-4 over 5= 4 then 1over 3

OpenStudy (anonymous):

final answer is v=13/3+4/5..so by l.c.m..we get 77/15!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome..!!